OpenStudy (anonymous):

Does (1/Sin^2x) = Csc^2x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright so if i have (1/Sin^2x) + (1/Cos^2x) = (Csc^2x)(Sec^2x) What would the proof look like?

OpenStudy (anonymous):

let me do it on paper real quick

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

no you have the wrong equation \[\frac{1}{\sin^2(x)} + \frac{1}{\cos^2(x)} = \frac{\cos^2(x) + \sin^2(x)}{\sin^2(x)\cos^2(x)}\] which can be simplified

OpenStudy (anonymous):

This is a problem on my trig homework though

OpenStudy (anonymous):

get common denominator on the left

OpenStudy (campbell_st):

which means its \[\frac{1}{\sin^2(x)\cos^2(x)} = \frac{1}{\sin^2(x)} \times \frac{1}{\cos^2(x)}\] you should be able to finish it from here

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Heres another proof

OpenStudy (anonymous):

Secx - Cosx = SinxTanx

OpenStudy (campbell_st):

well sec is 1/cos so you need a common denominator so the left hand side becomes \[\frac{1 - \cos^2(x)}{\cos(x)}\] make the obvious substitution to the numerator and then split it to sin(x) x sin(x)/cos(x) which give the solution

OpenStudy (anonymous):

But that gives Tanx x Tanx

OpenStudy (anonymous):

nevermind

OpenStudy (campbell_st):

no make the substitution and you get \[\frac{\sin^2(x)}{\cos(x)} = \frac{\sin(x) \times \sin(x)}{\cos(x)}\] hope this helps

OpenStudy (campbell_st):

then if you split the numerator you have \[\sin(x) \times \frac{\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

Another Question, ... Tanx + Cotx = SecxCscx

OpenStudy (campbell_st):

well tan = sin/cos and cot= cos/sin to you have sin/cos + cos/sin you'll need a common denominator... and cross multiply... then simplify

OpenStudy (anonymous):

So (Cos^2x + Sin^2x) / CosxSinx

OpenStudy (anonymous):

yes or you could simplify the top because (cos^2x+sin^2x)=1

OpenStudy (anonymous):

as long as it matches it doesnt really matter

OpenStudy (anonymous):

Oh okay thank you so much

OpenStudy (anonymous):

i gave you a medal gotta go bye